Potassium chloride 10%
Cherry syrup q.s. ad 480 mL
Sig. tablespoonful b.i.d.

How many milliequivalents of potassium chloride are represented in each prescribed dose?
Answer #1

Absolute valence = 1x1 = 1

Molecular weight of kCl = 74.5

mEq = (mg x valence) / mw

Or 1mEq = mw / valence

10% = 10g / 100 ml

1 tbsp = 15ml ——> (10g/100ml) x 15ml = 1.5g = 1500mg

mEq = (mg x valence) / mw = (1500mg x 1) / 74.5 = 20.13 mEq KCl